A telecommunications company has four satellites, two of which are sending a weak signal. If two are picked at random without replacement, find the probability that both are sending a weak signal. Express your answer as a fraction in lowest terms.


1
Expert's answer
2021-11-15T19:24:47-0500

"P=\frac{2}{4}*\frac{1}{3}=\frac{1}{6}."